# TempFlow-GRPO: When Timing Matters for GRPO in Flow Models

### v2

@@ A.4 Comparsion with DanceGRPO
Figure 12: (Left) Performance of TempFlow-GRPO on HPDv2 with HPSv2 as reward compared DanceGRPO. (Right) Comparsion on PickScore benchmark (PickScore reward, FLUX.1-dev 1024).
@@ A.9 Multi-Reward Optimization
We evaluate the performance of our method in multi-reward optimization scenarios. For this experiment, we employ FLUX.1-Dev as the base model and train at a high resolution of $\times1024$. We simultaneously optimize for two reward functions—HPSv3 and PickScore—with a weight ratio of 1:0.26. After 120 training steps, the results presented in Table 2 demonstrate our method’s robust capability for multi-objective optimization. Compared to Flow-GRPO, our approach achieves a substantial improvement of 0.82 on the HPSv3 metric while simultaneously maintaining consistent performance gains on PickScore. These results validate that our temporal reweighting mechanism effectively balances multiple objectives without sacrificing performance on individual metrics, highlighting the versatility of our approach in complex preference alignment scenarios where multiple criteria must be optimized concurrently.
@@ A.10 Visualization
1. 16-year-old teenager wearing a white bear-ear hat with a smirk on their face. 2. photo of well done salmon dinner, 8K, Global Illumination, Ray Tracing Reflections 3. A lemon with a McDonald’s hat. 4. An image of an emo with dark brown hair in a messy pixie cut, large entirely-black eyes, wearing black clothing. 5. The image is a mixed media collage with broken glass and torn paper elements, featuring intricate oil details and a canvas texture, in a contemporary art style. 6. An epic deep space photograph. Gigantic, monolithic letters forming the word ’TempFlow-’ hang silently in the void, their surfaces like ancient, cracked obsidian reflecting distant starfields. Far below, the letters ’GRPO’ are formed by a vast, tranquil nebula glowing with soft, ethereal light, like a cosmic ocean. The sense of scale is immense and humbling. Perfect, case-sensitive lettering. Moody, atmospheric, photorealistic, cinematic wide-angle shot, 4K UHD. 7. Kiwi fruit, mint leaves, ice cubes, background yellow, splashing water, soft box, back light, creative food photography, Art by Alberto Seveso, 8. Claymation of Futurama characters. 9. A group of four friends commemorating a ski trip in the snow. 10. an empty bench next to a busy street. photo of well done salmon dinner, 8K, Global Illumination, Ray Tracing Reflections 11. a 12 year old girl and her pet raccoon

### v3

@@ A.4 Comparsion with DanceGRPO
Figure 12: (Left) Performance of TempFlow-GRPO on HPDv2 with HPSv2 as reward compared DanceGRPO. (Right) Comparsion on PickScore benchmark (PickScore reward, SD3.5-M 1024).
@@ A.9 Multi-Reward Optimization
We evaluate the performance of our method in multi-reward optimization scenarios. For this experiment, we employ FLUX.1-dev as the base model and train at a high resolution of $\times1024$. We simultaneously optimize for two reward functions—HPSv3 and PickScore—with a weight ratio of 1:0.26. After 120 training steps, the results presented in Table 2 demonstrate our method’s robust capability for multi-objective optimization. Compared to Flow-GRPO, our approach achieves a substantial improvement of 0.82 on the HPSv3 metric while simultaneously maintaining consistent performance gains on PickScore. These results validate that our TempFlow-GRPO effectively balances multiple objectives without sacrificing performance on individual metrics, highlighting the versatility of our approach in complex preference alignment scenarios where multiple criteria must be optimized concurrently.
@@ A.10 Visualization
1. 16-year-old teenager wearing a white bear-ear hat with a smirk on their face. 2. photo of well done salmon dinner, 8K, Global Illumination, Ray Tracing Reflections 3. A lemon with a McDonald’s hat. 4. An image of an emo with dark brown hair in a messy pixie cut, large entirely-black eyes, wearing black clothing. 5. The image is a mixed media collage with broken glass and torn paper elements, featuring intricate oil details and a canvas texture, in a contemporary art style. 6. An epic deep space photograph. Gigantic, monolithic letters forming the word ’TempFlow-’ hang silently in the void, their surfaces like ancient, cracked obsidian reflecting distant starfields. Far below, the letters ’GRPO’ are formed by a vast, tranquil nebula glowing with soft, ethereal light, like a cosmic ocean. The sense of scale is immense and humbling. Perfect, case-sensitive lettering. Moody, atmospheric, photorealistic, cinematic wide-angle shot, 4K UHD. 7. Kiwi fruit, mint leaves, ice cubes, background yellow, splashing water, soft box, back light, creative food photography, Art by Alberto Seveso, 8. Claymation of Futurama characters. 9. A group of four friends commemorating a ski trip in the snow. 10. an empty bench next to a busy street. 11. a 12 year old girl and her pet raccoon